OpenStudy (anonymous):

How would I multiply this without a calculator? 2sqrt(2)/sqrt(17)*(sqrt(17)/sqrt(17))

OpenStudy (anonymous):

\[\frac{2\sqrt{2}}{\sqrt{17}} * \frac{\sqrt{17}}{\sqrt{17}}\]

OpenStudy (anonymous):

write with brackets....its not clear..

OpenStudy (ash2326):

We have \[\frac{2\sqrt 2}{\sqrt{17}}* \frac{\sqrt{17}}{\sqrt{17}}\]

OpenStudy (anonymous):

2sqrt(2)/sqrt(17) = 2sqrt34/ 17 (sqrt(17)/sqrt(17)) = 1 result is 2sqrt34/17

OpenStudy (anonymous):

So... \[\frac{2\sqrt 2}{\sqrt{17}}* \frac{\sqrt{17}}{\sqrt{17}} = \frac{2\sqrt{34}}{17}\] is that what your saying?

OpenStudy (ash2326):

Yeah, you're correct

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So you just multiply under the radical you don't multiply outside by the 2 also.

OpenStudy (anonymous):

OK Thanks... So Much. Radicals are my bane!

OpenStudy (anonymous):

yw